OpenStudy (anonymous):

let y=a/(bx^2). Describe how y will change. a.) double a, keeping b and x constant. b.) double x, keeping a and b constant.

OpenStudy (anonymous):

a

OpenStudy (john_es):

a) If you double a, then, \[y_{new}=2a/(bx^2)=2y\]so y duplicates its value. b) If you double x, then \[y_{new}=a/(b(2x)^2)=a/(4bx^2)=y/4\]so y gets divides by 4.